# State of practice: evaluating GPU performance of state vector and tensor network methods

### v1

@@ Abstract
The frontier of quantum computing (QC) simulation on classical hardware is quickly reaching the hard scalability limits for computational feasibility. Nonetheless, there is still a need to simulate large quantum systems classically, as the Noisy Intermediate Scale Quantum (NISQ) devices are yet to be considered fault tolerant and performant enough in terms of operations per second. Each of the two main exact simulation techniques, state vector and tensor network simulators, boasts specific limitations. The exponential memory requirement of state vector simulation, when compared to the qubit register sizes of currently available quantum computers, quickly saturates the capacity of the top HPC machines currently available. Tensor network contraction approaches, which encode quantum circuits into tensor networks and then contract them over an output bit string to obtain its probability amplitude, still fall short of the inherent complexity of finding an optimal contraction path, which maps to a max-cut problem on a dense mesh, a notably NP-hard problem.
This article aims at investigating the limits of current state-of-the-art simulation techniques on a test bench made of eight widely used quantum subroutines, each in 31 different configurations, with special emphasis on performance. We then correlate the performance measures of the simulators with the metrics that characterise the benchmark circuits, identifying the main reasons behind the observed performance trend. From our observations, given the structure of a quantum circuit and the number of qubits, we highlight how to select the best simulation strategy, obtaining a speedup of up to an order of magnitude.
@@ I Introduction
The quantum computing paradigm, since its theoretical inception by one of the forefathers of modern physics R. Feynman in 1982[14], served to extend the classical definition of computation in order to better describe the quantum properties of nature, under the hypothesis that an experiment, which is a purposefully built physical system, can be said to be performing computations under specific conditions[20]. Years later, many quantum algorithms would be proposed, able to exploit the characteristics of this new paradigm to achieve unprecedented speedups in unstructured search problems [17] or to efficiently solve hard mathematical problems, such as the prime factorization of large numbers [28], and others in the fields of optimisation, machine learning and chemistry.
@@
In the last decade, the progress in the various technologies used for building quantum devices has reached commercial applications, such as the cloud services proposed by IBM, D-Wave and Google, among others. These systems, however, are yet to be considered mature, as most still fail in preserving coherent quantum states, suffer cross-talk among their constituent elements and employ imperfect logical operators. Furthermore, supercondution-based devices, which make up a wide margin of the operational quantum devices as of the writing of this article, have been recently proven to be exceedingly susceptible to external radiation events [35, 37, 24], once again hindering their applicability.
@@
Meanwhile, the top performing supercomputers have surpassed the exascale number of floating operations per second [32]. Naturally, such an achievement pushes up the threshold for the complexity of quantum systems that can be classically simulated over a reasonable time span [19]. This opens up the possibility for some quantum algorithms to be efficiently converted into quantum-inspired classical algorithms. Moreover, the testing and validation of newly proposed quantum algorithms, and the accuracy measurement of real quantum devices’ outputs must be done through simulation, to double check results and ensure correct operation. With the increasingly higher optimisation of specialised software libraries over commodity hardware accelerators, such as graphic processing units (GPUs), simulation of small quantum algorithms is becoming more and more easy to accomplish [29, 30, 1, 5]. Such practices are, of course, still limited in memory and time by the exponential requirements of quantum simulation, but better exploitation of current classical computation resources may lead to efficient simulation of small quantum subroutines without needing to compensate for hardware-level noise and external radiation events, possibly reducing or annulling the queries made to cloud-based quantum computers for low qubit sized problems.
@@
The purpose of this article is to understand where the limit for efficient quantum simulation on classical hardware lies, emphasising the computational aspects, including the time and memory footprints of quantum algorithms. There are various classical simulation techniques for quantum algorithms other than state vector simulation and tensor network contractions, such as stabilisers theory or p-block simulation. Notably, however, stabiliser theory is limited to the application of a non-universal set of Clifford gates, whilst p-block simulation leverages approximations in the representation that lead to very limited entanglement. Being the only two approaches able to find an exact solution, the analysis we propose focuses on state vector simulation and tensor network contraction techniques.
@@
We prove that, by profiling quantum circuits with the approach presented in this paper, the simulation time can reach a speedup of up to one order of magnitude, especially for large quantum circuits. The proposed performance-circuit characteristics correlation is achieved by characterising a purposefully selected suite of well known quantum circuit subroutines according to objective metrics, and checking how those scale with respect to the size of the quantum circuit. All the circuits we consider are parameterisable over the number of qubits in the system, and some of them feature additional customisation parameters, such as layer repetition. Moreover, they have been selected as to have practical applicability in terms of exact simulation. These same circuits will be then simulated on top level hardware, using both state-vector and tensor network contraction approaches through NVIDIA’s cuQuantum library [29], highlighting which one boasts the better performance for each workload. Our work proposes a practical methodology to pick the most efficient simulation strategy according to a given set of static characteristics of the circuit.
@@
The rest of the paper is organized as follows: Section II provides a short summary of quantum computing as a whole, and it is followed by Section III, which introduces how classical algorithms for quantum simulation work. Section IV gives a definition of the metrics and of the quantum circuits considered for this study. Section V characterises the quantum circuits according to the aforementioned metrics, then presents performance results with respect to execution times and peak memory occupancy. Lastly, Section VI concludes the paper by expanding on the hereby presented work by opening new paths for investigation in future works.
@@ II-A Quantum Computing
Entanglement is believed to be the fundamental resource responsible for quantum speedup, although superposition plays an important role as well, since circuits with low entanglement have been proven to be easy to simulate [36].
@@ III-A State vector simulation
The state vector simulation’s complexity scales linearly in time with respect to the number of gates [12]. However, the memory footprint of the state vector and the number of vector-matrix multiplications performed increase exponentially with the number of qubits present in the system to be simulated.
[…]
To put that into perspective, it is possible to roughly estimate the number of atoms in the observable universe to be $10^{82}\approx 2^{270}$ [27]: this means that, if we were to to store a single amplitude value inside each of them to represent a state vector, we could only represent state vectors of systems with up to 270 qubits. Well known quantum algorithms need significantly more logical qubits [18], and this is without considering the cost in terms of classical computation time, which may add up to reach unfathomable time scales[3]. Overall, the state vector approach is generally convenient when simulating small quantum systems, as it produces a full description of the output wavefunction.
@@ III-B Tensor network simulation
All other gates use their standard matrix representation. The contraction of an edge corresponds to the multi-dimensional generalisation of the dot product between two tensors over a shared index. The measurement operators at the end of the quantum circuit are substituted by open indices. Whenever a full network contraction operation is performed, the output open indices are closed with the conjugate tensors of the basis states that represent a specific bit string. Doing so, followed by the contraction of the newly closed indices, produces the probability amplitude of the chosen bit string. Figure 4 provides graphical insight over this process.
@@
[…]
The memory occupancy of the tensor network grows linearly with respect to the number of quantum gates and qubits in the system. This approach moves the complexity of simulation to that of finding an optimal contraction path for the tensor network, which is known to be an NP-hard problem [15]. Efficient heuristics specialised for quantum circuit-derived tensor networks have been proposed [16], however there is no catch all heuristic for this kind of problem in general. If the contraction path is not optimal, it may lead to increased computation time, possibly making it less efficient than state vector simulation altogether. To the interested reader, we suggest some resources for tensor network theory by J. Biamonte [8, 7].
@@ IV Benchmarks and metrics
[…]
@@ IV-B1 Entanglement variance
QAOA
Quantum Approximate Optimisation Algorithm
$\frac{3}{2}PN(N-1)+2N$
$PN(N-1)$
[13]
Random
Google quantum supremacy circuit
$(1-k)(N(\lfloor N/2\rfloor+N\%2))+kN^{2}$
$kN(\lfloor N/2\rfloor)$
[3]
QPE
Quantum Phase estimation
$\frac{N(N-1)}{2}+2N-1+\lfloor\frac{(N-1)}{2}\rfloor$
$\frac{(N^{2}-N)}{2}+N-2+\lfloor\frac{(N-1)}{2}\rfloor$
[21]
QFT
Quantum Fourier transform
$\frac{1}{2}N(N+1)+\lfloor N/2\rfloor$
$\frac{1}{2}(N^{2}-N)+\lfloor N/2\rfloor$
[10]
VQE
Variational Quantum Eigensolver
$L(5N-1)+N$
$L(N-1)$
[26]
Hamiltonian sim.
One-dimensional Hamiltonian time evolution
$3T(2N-1)$
$T(N-1)$
[31]
Hidden Shift
Find the shift $s$ such that $g(x)=f(x+s)$
$3N+2M+\lfloor N/2\rfloor$
$\lfloor N/2\rfloor$
[34]
Bernstein-Vazirani
Hidden bit string extraction
$2N+M$
$M$
[6]
Figure 5: The scaling of the number of total gates of the circuits considered in this paper, from size 2 to size 32. The Y-axis is in logarithmic scale. The controlling variables are set as $P=1$, $k=0.5$, $M=\lfloor kN\rfloor$, $L=1$, $T=1$. The interactive version of this plot is available online [33].
@@ V Results
We will first characterise the quantum circuits according to the metrics introduced in Section IV, then we will relate such metrics with the execution time and memory occupancy metrics of each quantum circuit. The library used for running the simulations is the NVIDIA cuQuantum library [29], which has been cross-compiled with the QSim simulator library [30]. This gives us access to three different GPU accelerated simulation backends:
@@
All experiments have been run on a system equipped with an AMD EPYC 7643 48-Core Processor, 128 GB of RAM and an NVIDIA A100 80 GB GPU.
@@ V-A Quantum Circuit Properties
Figure 8: At the top, the execution times for the VQE circuit (left) and the Random circuit (right). All data has been collected by doing 3 warmup runs and then averaging the results of 10 additional runs. The tensor network time is the sum of the pathfinding time, done in CPU, and the contraction time, done in GPU. At the bottom, the network structure derived from the same circuits of size 32 qubits, elided of the tensors with degree $d=2$. Blue nodes represent input tensors, while red nodes represent output sensors, node size represents the order of the tensor.
@@ V-B Simulation Performance
In Figure 8, we can see a side by side comparison of the execution times of the VQE and the Random circuits, together with the tensor network structure derived from the circuit, after the removal of all order-2 tensors, for readability’s sake. The time performance data has been collected, for each circuit configuration, as the average time over 10 runs after having performed 3 warmup runs. Thus we collectively performed $3244$ quantum circuit simulations. Code and execution time data relative to these other circuits is available at [2]. We notice how the qsim-cuda simulator has the overall fastest performance for circuits with size of 13 qubits or less, after which the performance of the simulator degrades significantly. This is mainly due to the fact that the shared memory size of the NVIDIA A100, the device we used for simulations, is only configurable up to 164 KB: the state vector of a 13 qubits system, at complex-64 single precision, occupies 64 KB. The performance for circuits of size equal to 14 still holds up, as the state vector needs 128 KB to be stored. However, as soon as the size increases, the L1 cache is not large enough to fit the whole state vector and the simulation gets hindered by memory transfers. From that point onward, the execution time for the qsim-cuda scales exponentially with the system size. Quantum circuits with more than 14 qubits perform better on either of the other two simulators. Notably, the qsim-cusv simulator starts to lose performance for system sizes larger than 22 qubits. This is once again due to the properties of the A100 GPU, which has an L2 cache size of 40 MB, whilst the state vector size of a 22 qubit system is 32 MB. As soon as we increase the system size, the memory requirement doubles and exceeds the cache, forcing additional memory transfers that make the time performance once again scale exponentially in time.
In the case of the Hamiltonian simulation, VQE and QAOA circuits, the tensor network outperforms both state vector simulators. This is a consequence of the fact that these circuits are mainly composed of single qubit gates, corresponding to an entanglement ratio score lower than $0.5$, and have a well distributed amount of entanglement across the system, meaning an entanglement variance score lower than $0.2$. This leads to having small intermediate tensors during the contraction step, that eventually get merged together into larger tensors right towards the end of the process. As soon as one of these two metrics raises up too much, the current pathfinding algorithm [16] cannot find a good enough contraction path.
@@
[…]
Moreover, once all the order-2 tensors have been contracted, the resulting topology resembles that of a matrix product state (MPS), which is optimal for contraction. This is an important information, as the contractions of order-2 tensors are easy to perform and do not increase the intermediate tensor order during the contraction process. The circuit with the worst performance for the cutn simulator is the Random circuit, which is the only problem where performance degrades by more than an order of magnitude, mainly due to its lack of an internal structure, something which can be clearly noticed on the right of Figure 8. In all other benchmark circuits, the performance in terms of time is comparable to that of a state vector simulator, whilst keeping the reduced memory footprint of the tensor network representation.
@@
Figure 10: The decision diagram to be used for picking the most fitting simulation backend given an arbitrary quantum circuit.
V-C Lessons Learnt

### v2

@@ Abstract
The frontier of quantum computing (QC) simulation on classical hardware is quickly reaching the hard scalability limits for computational feasibility. Nonetheless, there is still a need to simulate large quantum systems classically, as the Noisy Intermediate Scale Quantum (NISQ) devices are yet to be considered fault tolerant and performant enough in terms of operations per second. Each of the two main exact simulation techniques, state vector and tensor network simulators, boasts specific limitations.
This article investigates the limits of current state-of-the-art simulation techniques on a test bench made of eight widely used quantum subroutines, each in different configurations, with a special emphasis on performance. We perform both single process and distributed scaleability experiments on a supercomputer. We correlate the performance measures from such experiments with the metrics that characterise the benchmark circuits, identifying the main reasons behind the observed performance trends. Specifically, we perform distributed sliced tensor contractions, and we analyse the impact of pathfinding quality on contraction time, correlating both results with topological circuit characteristics. From our observations, given the structure of a quantum circuit and the number of qubits, we highlight how to select the best simulation strategy, demonstrating how preventive circuit analysis can guide and improve simulation performance by more than an order of magnitude.
@@ I Introduction
The quantum computing paradigm, since its theoretical inception by one of the forefathers of modern physics R. Feynman in 1982[1], served to extend the classical definition of computation in order to better describe the quantum properties of nature, under the hypothesis that an experiment, which is a purposefully built physical system, can be said to be performing computations under specific conditions[2]. Years later, many quantum algorithms would be proposed, able to exploit the characteristics of this new paradigm to achieve unprecedented speedups in unstructured search problems [3] or to efficiently solve hard mathematical problems, such as prime factorization of large numbers [4], and other more recent applications in the fields of optimisation [5], machine learning [6] and chemistry [7].
@@
In the last decade, progress in the various technologies used for building quantum devices has reached commercial applications, such as the cloud services proposed by IBM, D-Wave and Google, among others. These systems, however, are yet to be considered mature, as most still fail in preserving coherent quantum states, suffer cross-talk among their constituent elements and employ imperfect logical operators. Furthermore, supercondution-based devices, which make up a wide margin of the operational quantum devices as of the writing of this article, have been recently proven to be exceedingly susceptible to external radiation events [8, 9, 10], once again hindering their applicability.
@@
Meanwhile, the top performing supercomputers have surpassed the exascale number of floating operations per second [11]. Naturally, such an achievement pushes up the threshold for the complexity of quantum systems that can be classically simulated over a reasonable time span [12]. This opens up the possibility for some quantum algorithms to be efficiently converted into quantum-inspired classical algorithms. Moreover, the testing and validation of newly proposed quantum algorithms, and the accuracy measurement of real quantum devices’ outputs must be done through simulation, to double check results and ensure correct operation. With the increasingly higher optimisation of specialised software libraries over commodity hardware accelerators, such as graphic processing units (GPUs), the rapid simulation of small quantum algorithms is becoming more and more easy to accomplish [13, 14, 15, 16]. Such practices are, of course, still limited in memory and time by the exponential requirements of quantum simulation, but better exploitation of current classical computation resources may lead to efficient simulation of small quantum subroutines without needing to compensate for hardware-level noise and external radiation events, possibly reducing or annulling the queries made to cloud-based quantum computers for low qubit sized problems.
@@
The purpose of this article is to understand where the limit for efficient quantum simulation on classical hardware lies, emphasising the computational aspects, such as distributed performance, scaleability, time and memory footprints of quantum algorithms, with the objective to find quantum circuit features that correlate to simulation performance. There are various classical simulation techniques for quantum algorithms other than state vector simulation and tensor network contractions, such as stabilisers theory or p-block simulation. Notably, however, stabiliser theory is limited to the application of a non-universal set of Clifford gates, whilst p-block simulation leverages approximations in the representation that lead to very limited entanglement. Being the only two approaches able to find an exact solution, the analysis we propose focuses on state vector simulation and tensor network contraction techniques.
The questions seek an answer for are: what is the performance of state of the art quantum simulation methods? Which topological features of a quantum circuit correlate to simulation performance, and which simulation approach is more suitable? Are there limitations to distributed quantum simulation, and can we predict them?
@@
We prove that, by profiling quantum circuits with the approach presented in this paper, the simulation time can reach a speedup of up to one order of magnitude, especially for large quantum circuits, on a single GPU. Furthermore, we report the results from distributed tensor contraction simulations, highlighting speedups of more than 364×364$\times$364 × with respect to single GPU performance, and the we trace the impact of pathfinding quality on the contraction performance, obtaining speedups of up to 4.79×4.79$\times$4.79 × through tuning. The proposed circuit metrics to performance correlation is achieved by characterising a purposefully selected suite of well known quantum circuit subroutines according to objective metrics, and checking how those scale with respect to the size of the quantum circuit. All the circuits we consider are parameterisable over the number of qubits in the system, and some of them feature additional customisation parameters, such as layer repetition. Moreover, they have been selected as to have practical applicability in terms of exact simulation. These same circuits have been simulated on CINECA’s Leonardo supercomputer, the $7^{th}$ supercomputer in the Top500 list, using both state-vector and tensor network contraction methods through NVIDIA’s cuQuantum library [13], highlighting which one boasts the better performance for each workload. Our work proposes a practical methodology to pick the most efficient simulation strategy according to a given set of static characteristics of the circuit.
@@
The rest of the paper is organized as follows: Section II provides a short summary of quantum computing as a whole, and it is followed by Section III, which introduces how classical algorithms for quantum simulation work. Section IV gives a definition of the metrics and of the quantum circuits considered for this study. Section V characterises the quantum circuits according to the aforementioned metrics, then presents performance results with respect to execution times and peak memory occupancy, scaling of distributed tensor network contractions and impact of pathfinding resources on tensor network contraction times. Lastly, Section VI concludes the paper by expanding on the hereby presented work by opening new paths for investigation in future works.
@@ II-A Quantum Computing
Entanglement is believed to be the fundamental resource responsible for quantum speedup, although superposition plays an important role as well, since circuits with low entanglement have been proven to be trivial to simulate [19].
@@ III-A State vector simulation
The state vector simulation’s complexity scales linearly in time with respect to the number of gates [21]. However, the memory footprint of the state vector and the number of vector-matrix multiplications performed increase exponentially with the number of qubits present in the system to be simulated, so this approach is not scaleable indefinitely. To put that into perspective, it is possible to roughly estimate the number of atoms in the observable universe to be $10^{82}\approx 2^{270}$ [24]: this means that, if we were to to store a single amplitude value inside each of them to represent a state vector, we could only represent state vectors of systems with up to 270 qubits. Well known quantum algorithms need significantly more logical qubits [25], and this is without considering the cost in terms of classical computation time, which may add up to reach unfathomable time scales[26]. Overall, the state vector approach is generally convenient when simulating small quantum systems, as it produces a full description of the output wavefunction.
@@ III-B Tensor network simulation
All other gates use their standard matrix representation. The contraction of an edge corresponds to the multi-dimensional generalisation of the dot product between two tensors over a shared index. The measurement operators at the end of the quantum circuit are substituted by open indices. Whenever a full network contraction operation is performed, the output open indices are closed with the conjugate tensors of the basis states that encode a specific bit string. Doing so, followed by the contraction of the newly closed indices, produces the probability amplitude of the chosen bit string. Figure 4 provides graphical insight over this process.
@@
Observation I Tensor network contractions can reconstruct the full state vector, and such process can be trivially parallelised over different bitstrings.
The memory occupancy of the tensor network grows linearly with respect to the number of quantum gates and qubits in the system. This approach moves the complexity of simulation to that of finding an optimal contraction path for the tensor network, which is known to be an NP-hard problem [27]. Efficient heuristics specialised for quantum circuit-derived tensor networks have been proposed, however there is no catch all solution for this kind of problem. The pathfinding algorithm used in this work [22], despite representing the state-of-the-art for this class of problems, only strives to optimise having the lowest possible amount of floating-point operations across the whole contraction process, which does not prevent the formation of large intermediate tensors, something that inevitably reduces contraction performance. Besides, as it will be discussed in Section V-D, the optimiser may easily be locked in a local minimum in some problems, whereas other problems feature smoother landscapes in terms of pathfinding complexity. If the contraction path is not optimal, it may lead to increased computation time, possibly making it less efficient than state vector simulation altogether. To the interested reader, we suggest some resources for tensor network theory by J. Biamonte [28, 29].
@@ IV Benchmarks and metrics
Observation II Not all quantum circuits generally used for benchmarking are computationally representative in a classical simulation environment.
@@ IV-B1 Entanglement variance
QAOA Quantum Approximate Optimisation Algorithm $\frac{3}{2}PN(N-1)+2N$ $PN(N-1)$ [5]
Random Google quantum supremacy circuit $(1-k)(N(\lfloor N/2\rfloor+N\%2))+kN^{2}$ $kN(\lfloor N/2\rfloor)$ [26]
QPE Quantum Phase estimation $\frac{N(N-1)}{2}+2N-1+\lfloor\frac{(N-1)}{2}\rfloor$ $\frac{(N^{2}-N)}{2}+N-2+\lfloor\frac{(N-1)}{2}\rfloor$ [32]
QFT Quantum Fourier transform $\frac{1}{2}N(N+1)+\lfloor N/2\rfloor$ $\frac{1}{2}(N^{2}-N)+\lfloor N/2\rfloor$ [33]
VQE Variational Quantum Eigensolver $L(5N-1)+N$ $L(N-1)$ [6]
Hamiltonian sim. One-dimensional Hamiltonian time evolution $3T(2N-1)$ $T(N-1)$ [30]
Hidden Shift Find the shift $s$ such that $g(x)=f(x+s)$ $3N+2M+\lfloor N/2\rfloor$ $\lfloor N/2\rfloor$ [34]
Bernstein-Vazirani Hidden bit string extraction $2N+M$ $M$ [35]
Figure 5: The scaling of the number of total gates of the circuits considered in this paper, from size 2 to size 32. The Y-axis is in logarithmic scale. The controlling variables are set as $P=1$, $k=0.5$, $M=\lfloor kN\rfloor$, $L=1$, $T=1$.
@@ V Results
In order to find an answer to the research questions defined in Section I, we will start by characterising the quantum circuits according to the metrics introduced in Section IV, then we will perform various simulations, with the objective to relate the metrics with execution time, memory occupancy, and in the specific case of tensor networks, distributed sliced contraction performance and pathfinding efficacy. The library used for running the simulations is the NVIDIA cuQuantum library (version v24.03)[13], adapted to the specific needs of our analysis. This gives us access to three different GPU accelerated simulation backends:
@@
All experiments have been run on CINECA’s Leonardo supercomputer. Apart from the distributed experiments, all other experiments have been run on a single node, using 8 cores of an Intel Xeon Platinum 8358 CPU, 128 GB of RAM and one NVIDIA Ampere A100 64 GB GPU. Simulations have all been limited to a problem size of 32 qubits, as that is the largest statevector that can be represented in a single available GPU, although larger tensor networks could indeed be simulated. Given the high computational cost of state vector and tensor network methods, CPU-based simulation algorithms will not be considered for this analysis, since it would not provide any meaningful comparison in terms of performance.
@@
Figure 7: The execution times for the VQE circuit (left) and the Random circuit (right). All data has been collected by doing 3 warmup runs and then averaging the results of 10 additional runs. The tensor network time is the sum of the pathfinding time, done in CPU, and the contraction time, done in GPU.
@@ V-B Single GPU simulation performance
In Figure 7, we can see a side by side comparison of the execution times of the VQE and the Random circuits. The time performance data has been collected, for each circuit configuration, as the average time over 10 runs after having performed 3 warmup runs. Thus we collectively performed $3244$ quantum circuit simulations for this experiment. Code and execution time data relative to these other circuits is available at [37]. We notice how the qsim-cuda simulator has the overall fastest performance for circuits with size of 13 qubits or less, after which the performance of the simulator degrades significantly. This is mainly due to the fact that the shared memory size of the NVIDIA A100, the device we used for simulations, is only configurable up to 164 KB: the state vector of a 13 qubits system, at complex-64 single precision, occupies 64 KB. The performance for circuits of size equal to 14 still holds up, as the state vector needs 128 KB to be stored. However, as soon as the size increases, the L1 cache is not large enough to fit the whole state vector and the simulation gets hindered by memory transfers. From that point onward, the execution time for the qsim-cuda scales exponentially with the system size. Quantum circuits with more than 14 qubits perform better on either of the other two simulators. Notably, the qsim-cusv simulator starts to lose performance for system sizes larger than 22 qubits. This is once again due to the properties of the A100 GPU, which has an L2 cache size of 40 MB, whilst the state vector size of a 22 qubit system is 32 MB. As soon as we increase the system size, the memory requirement doubles and exceeds the cache, forcing additional memory transfers that make the time performance once again scale exponentially in time.
@@
In the case of the Hamiltonian simulation, VQE and QAOA circuits, the tensor network outperforms both state vector simulators. This is a consequence of the fact that these circuits are mainly composed of single qubit gates, corresponding to an entanglement ratio score lower than $0.5$, and have a well distributed amount of entanglement across the system, meaning an entanglement variance score lower than $0.2$. This leads to having small intermediate tensors during the contraction step, that eventually get merged together into larger tensors right towards the end of the process. As soon as one of these two metrics raises up too much, the pathfinding algorithm we employed [22] struggles find an optimal contraction path. Moreover, once all the order-2 tensors have been contracted, the resulting topology resembles that of a matrix product state (MPS), which is optimal for contraction. This is an important information, as the contractions of order-2 tensors are easy to perform and do not increase the intermediate tensor order during the contraction process.
Observation III Unstructured and unbalanced tensor networks give rise to large intermediate tensors during contraction in function of their qubit size and program communication metric, hindering performance.
The circuit with the worst performance for the cutn simulator is the Random circuit, which is the only problem where performance degrades by more than an order of magnitude, mainly due to its lack of an internal structure, something which can be clearly noticed on the right of Figure 7. In all other benchmark circuits, the performance in terms of time is comparable to that of a state vector simulator, whilst keeping the reduced memory footprint of the tensor network representation.
@@
V-C Distributed sliced tensor contraction performance
In order to understand how the performance of tensor network contraction scales as we increase the number of GPUs, we designed a strong scaling experiment. We implemented a distributed version of the tensor network contraction algorithm, by leveraging the cuTensorNet library, MPI and NCCL, and ran scaling experiments for all of the circuits in the benchmark at size 32 qubits, apart from the Random circuit which was limited at size 28 qubits, by using an increasing number of GPUs and compute nodes on the Leonardo Supercomputer, with the objective to test the efficacy of tensor network slicing in improving contraction efficacy. The algorithm starts by spawning one MPI process for each available GPU, and first performs a distributed pathfinding on the whole network. The best path, selected according to the lowest FLOPs count, is broadcast to all other MPI processes thorugh the MPI communicator. The tensor network is then sliced in a number of sub-networks equal to the number of MPI processes, in order to provide to each MPI process, and thus each GPU, a comparable amount of FLOPS to be performed. Each GPU contracts its own sub-network, and all the partial results are reduced with a sum operation through the NCCL communicator, that yields the final amplitude result.
Figure 10: Strong scaling of distributed sliced tensor network contraction performance on circuits of size 32 qubits, with the exclusion of the Random circuit at size 28 qubits. The number of GPUs (# GPUs) also corresponds the number of MPI processes. Each point represents the mean of 30 datapoints.
Figure 11: Pathfinding time (orange) and contract time (blue) for tensor networks derived from circuits of size 32 qubits. The number of samples performed by the pathfinder are executed sequentially in order to extrapolate the total time required for this computation, whilst contractions are performed using a single GPU. Notice different y-axis limits for each circuit in the benchmark. Vertical bars represent the 90th percentile interval.
In Figure 10, we present the strong scaling results of this experiment, from using 1 GPU going up to 256 GPUs. The node configuration on Leonardo includes 4 GPUs per node, and each node is interconnected with an Nvidia Mellanox network, reaching up to 200 Gbit/s node to node transfer rates. Given the size of the partial results being reduced, network bandwidth does not act as a computational bottelenck. We can see from the strong scaling results that not all the quantum circuits selected in the benchmark exhibit large performance gains, particularly the Hidden Shift, VQE, Bernstein-Vazirani and Hamiltonian simulation. These circuits lack complexity in the structure of their circuit-derived tensor networks, as reported by their asymptotically decreasing program communication metric in relation to the size of the quantum circuit. The QAOA sees a noticeable improvement in terms of contraction performance, steadily lowering the contraction time from 81 ms on a single GPU to as low as 19 ms on 16 GPUs, a speedup of more than 4.2×4.2$\times$4.2 ×. Likewise, the Random circuit’s contraction time goes from a mean of $1.89$ s when using a single GPU to about 120 ms in the case with 256 GPUs, a speedup of about 15×15$\times$15 ×. The largest improvements in the contraction times can be obtained on the QFT and QPE quantum circuits, which drop more than one order of magnitude in contraction time when going from running on one GPU to 16 GPUs: we respectively measure a speedup of more than 294×294$\times$294 × on the QFT circuit and a speedup of more than 364×364$\times$364 × on the QPE circuit, in lieu of just increasing by 256×256$\times$256 × the computational resources available.
Observation IV Quantum circuits with large program communication scores benefit the most from sliced distributed contraction, reaching superlinear speedups with respect to a linear increase in computational power.
Albeit some of the benchmark circuits considered have been found to be trivially solvable even by a single GPU, we demonstrated how specific descriptive metrics of a quantum circuit, namely the program communication metric, can let us foresee the presence or absence of a computational gain with a distributed sliced tensor network contraction.
On the topic of evaluating distributed sliced performance, it is still unclear how to measure weak scaling performance of different quantum circuits, that is measuring the performance of a problem when scaling equally the problem’s complexity and the available computing resources. The problems considered in this benchmark are parameterised by the number of qubits, which does not provide direct control over the problem’s contraction complexity, which mainly depends on the treewidth of the tensor network [22]. It could be possible to develop a synthetic parameterisable quantum circuit that grows in terms of the circuit derived tensor network’s treewidth. This would most probably end up being a variation of a Random circuit, which however holds no meaning in terms of problem solution.
Observation V The complexity of contracting quantum circuit derived tensor networks does not scale with the problem size, i.e. the number of qubits. As such, specialised synthetic benchmarks are needed to measure the weak scaling of tensor network contraction.
For the sake of this article, we can safely predict a correlation with the strong scaling capacity of the cuTensorNet library in relation to the program communication metric of a quantum circuit. For quantum circuits with program communication scores of one can efficiently leverage large multi-GPU acceleration. On the opposite case, when the program communication approaches zero, one GPU can suffice, and no advantage is gained from increasing computational resources.
V-D Pathfinding impact on tensor contraction performance
There is a need to classify quantum circuit derived tensor networks according to their pathfinding complexity. Specifically, we are interested in predicting which circuits can be contracted with higher efficiency in correlation to an increase in the resources available to the pathfinding algorithm. We investigate how a variation in the resources available to the cuTensorNet pathfinding algorithm, namely the number of samples performed, impacts on the contraction time of the quantum circuit derived tensor networks in the benchmark. We measure the total time required by the pathfinding algorithm to sequentially compute all of the samples by enforcing a single thread. Although this search may be easily distributed, the total computation time required by the pathfinder is still the same. The size of each quantum circuit is kept fixed at 32 qubits, apart form the random circuit of size 28 qubits, due to representability reasons.
In Figure 11, we plot the variation in the pathfinding and contraction time with respect to an increase in the number of samples, and find three categorisations for the quantum circuits used in this benchmark: pathfinding-bound problems, contraction-bound problems and unbounded problems. Pathfinding-bound problems include the VQE and Hamiltonian simulation circuits, since the complexity landscape of their possible contraction paths is mostly flat. This means that such circuits provide no advantage in terms of contraction time when assigning additional resources to the pathfinding algorithm. Contraction-bound problems include the QAOA, Hidden Shift, Bernstein-Vazirani and Random circuits, and they are characterised by having a non-trivial pathfinding complexity landscape, but trivial contraction complexity. This means that providing additional resources to the pathfinder indeed gives rise to better solutions in terms of total FLOPS in contraction, but the overall difference in terms of contraction efficiency is unnoticeable. Unbounded problems include The QFT and QPE problems, since they show an direct correlation between the amount of resources provided to the pathfinding algorithm and a reduction in the contraction time. In fact, as the number of samples increases, in both cases we can measure a speedup in terms of contraction time of 1.9×1.9$\times$1.9 × on the QPE and of 4.79×4.79$\times$4.79 × on the QFT by increasing the pathfinding time by about 29×29$\times$29 ×. From these observations it follows that quantum circuits characterised by an entanglement ratio metric that converges to one map to more complex pathfinding problems that can efficiently leverage additional pathfinding resources, whilst if the same metric saturates to values lower than $0.4$, the problem is bounded, either in pathfinding or contraction.
Observation VI Quantum circuit derived tensor networks can be classified in pathfinding-bound problems, contract-bound problems, and unbounded problems. Only the latter show an anticorrelation between pathfinding time and contraction time.
Although this result might point towards diminishing returns, it must be noted that the pathfinding time was purposefully measured in a sequential manner on a single thread, but each sample is indeed independent from the other, as such using one independent thread and core per sample is going to keep the real-time pathfinding time constant. Moreover, in terms of real-world tensor contraction, contractions are performed with a batched approach, where the contraction path is computed only once, but the actual contraction may be performed hundreds or thousands of times, thus outweighing the steep inital pathfinding cost. Given the intense computational cost of pathfinding and its reliance on performing numerous samples, an open question remains on whether this algorithm can be further accelerated using GPUs in order to increase the number of concurrent samples while still fitting a low time bound.
V-E Lessons learnt